OpenStudy (anonymous):

How do I solve the following equations using the substitution method? 20x-5y=20 and y=4x+4

OpenStudy (mrnood):

20x-5y=20 y=4x+4 The second equation gives you a value for y 'in terms of x' so 'substitute' 4x+4 in th eFIRST equation wherever there is a y This will give you a new equation with only x in it Write that equation here.....

OpenStudy (anonymous):

20(4x+4)-5y=20

OpenStudy (mrnood):

NO - you must substitute wherever you saw a y, not an x

OpenStudy (anonymous):

20x-5(4x+4)=20

OpenStudy (mrnood):

That's right - now expand the bracket, and collect the x terms together

OpenStudy (anonymous):

20x-20x+20=20

OpenStudy (mrnood):

hmmm - I see a problem

OpenStudy (anonymous):

they cross each other out?

OpenStudy (mrnood):

You have not expanded the bracket properly

OpenStudy (anonymous):

I am not understanding. I took the y and replaced it with the 4x+4 then multiplied by the 5 from the 5y?

OpenStudy (mrnood):

20x-5y=20 y=4x+4 If you rearrange the first 5y=20x-20 If you multiply the second by 5 5y = 20x +20 These equations cannot both be true - I think there is an error in your question....

OpenStudy (anonymous):

So there is no solution?

OpenStudy (mrnood):

They are parallel lines, therefore do not have a solution. Have you typed the question properly?

OpenStudy (anonymous):

yes I confirmed the question is correct but one of the options on my assignment is that there is no solution or that there is infinitely many solutions